# Generalized Finite-time Optimal Control Framework in Stochastic Thermodynamics

### v2

@@ 4.1 Multiple control parameters
where eq.˜23 is represented in Einstein’s summation convention, which signifies the summation over the contracted index. $\vec{\lambda}$ is the vector of control parameters. The Hessian of $\mathcal{L}^{*}$ (mass) gives the metric tensor $g_{ij}$,
@@
The MinAP implies solving a variational optimization problem, $\Sigma_{qs}^{*}=\inf(\Sigma_{drv}^{qs})=\int_{0}^{\tau}dt\>\mathcal{L}_{drv}^{*}$. The geodesic equation for the control parameter that minimizes $\Sigma_{drv}^{qs}$ eq.˜23 reads,
@@
$\begin{split}\Gamma_{jk}^{i}=\frac{1}{2}g^{im}\left(\frac{\partial g_{mj}}{\partial{\lambda}^{k}}+\frac{\partial g_{mk}}{\partial\lambda^{j}}-\frac{\partial g_{jk}}{\partial\lambda^{l}}\right)\end{split}$ (26)
Equations˜23 and 25 are equivalent to eqs.˜5 and 6 for the multi-parameter slow-driving optimal control problem.
@@
Using eq.˜27, the optimal slow driving EP reads:
@@
The finite-time optimal control framework developed in section˜3 is based on the MinAP combined with the existence of the slow-driven geodesic $\mathcal{G}(\{\lambda\})$ and incorporating the possibility of kinks. Hence, multi-parameter finite-time optimal control protocol is rather trivial, provided that a slow driving geodesic eq.˜27 exists and is computed. The corresponding finite-time multi-parameter optimal protocol reads,
@@
Equations˜27, 28, 29 and 30 are a generalization of eqs.˜7, 8, 18 and 21, respectively, for the multi-parameter generalized optimal control problem.
@@ 4.2 Numerical computation of geodesic
We have focused on the exact analytical formulation of the GFTOC framework. However, analytically computing the exact geodesic $\mathcal{G}(\{\lambda\})$ is not always feasible for more sophisticated problems. To this end, a numerical framework is required to compute the geodesic/optimal protocols. Here, we highlight two different numerical algorithms that are implemented to calculate the geodesic. The first algorithm is based on the MinAP, where the geodesic $\mathcal{G}(\{\lambda\})$ is obtained by the minimization of the Lagrangian eq.˜23, and is extended to return the finite-time optimal protocol $\mathcal{G}_{\tau}(\{\lambda\})$, it reads,
@@
Input: $A_{i},A_{f},\tau,g_{ij}$, eqs.˜23, 29 and 30
1:Compute the geodesic $\mathcal{G}(A_{\alpha})$ obtained by the recursive minimization of eq.˜23, and that connects $\mathcal{G}(A_{i})$ and $\mathcal{G}(A_{f})$.
2:Calculate the finite-time optimal protocol $\mathcal{G}_{\tau}(t)$ and EP cost $\Sigma_{\tau}^{*}$ using eqs.˜29 and 30, respectively.
@@
Note that, Algorithm 1 is modified to directly compute $\mathcal{G}_{\tau}(\{\lambda\})$ instead of $\mathcal{G}(\{\lambda\})$, if the thermodynamic cost for kinks is assigned under the action minimization problem and the endpoint discontinuity in the protocol space is allowed by relaxing the boundary condition constraint. Algorithm 1 extends to the class of numerical optimization methods developed in Ref.[52, 53, 54, 24, 55, 56, 57, 58, 59] and related to machine learning techniques Ref.[60, 61, 62]. However, none of the previous numerical algorithms incorporates kinks as a result of their failure to assign a thermodynamic cost to kinks. To our knowledge, Algorithm 1 is the novel formulation that numerically computes finite-time optimal protocols with kinks and computes the associated finite-time driving EP cost.
A second alternative algorithm computes the finite-time optimal protocols by numerically solving the geodesic eq.˜25 as a boundary value problem. The algorithm for it reads:
@@
Input: $A_{i},A_{f},\tau,g_{ij},\Gamma_{ij}^{k}$, eq.˜25 and eq.˜29
1:Compute the geodesic $\mathcal{G}(A_{\alpha})$ by numerically solving eq.˜25 under the boundary value constraint $A_{i}$ and $A_{f}$ at $t=0$ and $t=\tau$.
2:Calculate the finite-time optimal protocol $\mathcal{G}_{\tau}(t)$ and EP cost $\Sigma_{\tau}^{*}$ using eqs.˜29 and 30, respectively.
@@
In contrast to Algorithm 1, Algorithm 2 can only compute $\mathcal{G}(A_{\alpha})$, and requires eq.˜29 to compute $\mathcal{G}_{\tau}(A_{\alpha})$. The numerical approach elaborated in this section broadens the applicability of the GFTOC when an exact analytical solution of the geodesic is not available.
@@ 5 Applications of GFTOC framework
5.1 Linear optimal control
Linear irreversible thermodynamics utilizes the most fundamental quadratic dissipation function $\mathcal{L}^{*}=kA_{\alpha}^{2}$ for the EPR [72, 73]. This leads to $\mathcal{L}_{drv}^{*}=k\dot{A}_{\alpha}^{2}$ with the mass $\partial_{A_{\alpha}}^{2}\mathcal{L}^{*}=2k$. Fixing $k=1/2$ implies a unit mass. Solving $\dot{A}_{\alpha}=v_{qs}$ gives $\mathcal{G}^{lin}(A_{\alpha})=A_{\alpha}$. Thus, the geodesic is a linear interpolation between $A_{\alpha}^{i}$ and $A_{\alpha}^{j}$ in $A_{\alpha}$ space, and the slow driving optimal EP and the corresponding optimal protocol read,
The linear optimal control formulated here is a trivial example of Euclidean geometry with a constant mass.
@@ 5.3 Stochastic particle in a harmonic trap and
A stochastic particle in a harmonic trap potential models an important equivalence class that describes a large class of physical systems, and hence has been the testbed of novel theoretical developments in ST. The stochastic particle satisfies the Boltzmann distribution, $P\propto e^{-k(x-m)^{2}}$, with the trap stiffness $k$ and centre $m$. In this case, these two relevant control parameters fully capture the statistical properties of the system. We consider a finite-time optimal control problem for the time-dependent change of the trap stiffness $k(t)$ and centre $m(t)$ [7, 17]. The inverse of the trap stiffness quantifies the covariance $(C)$ of the Gaussian distribution for the particle position, $C\propto 1/k$ [7]. Hence, by definition, the finite-time optimal control formulation for changing the centre and stiffness of the harmonic trap is equivalent to the optimal transport theory to change the mean and covariance of the Gaussian distribution for the position of the particles [39]. However, the results obtained from different methods do not match [7, 39, 51]. We resolve these discrepancies using the GFTOC framework.
We consider an optimal control problem for changing the trap stiffness and centre from $(m_{i},k_{i})$ to $(m_{f},k_{f})$. The mass for the centre and the stiffness of the trap are $g_{mm}=1$ and $g_{kk}=1/k^{3}$ [17]. This implies the geodesic function $\mathcal{G}(m)=m$ and $\mathcal{G}(k)=k^{-1/2}$. Using $C\propto 1/k$, we recover the geodesic function $\mathcal{G}(C)=C^{1/2}$ for the covariance. Hence, the minimum slow driving EP and the corresponding optimal protocols for the mean and variance read:
$\Sigma_{qs}^{*}=\frac{1}{2\tau}\left[(m_{f}-m_{i})^{2}+\left(\sqrt{C_{f}}-\sqrt{C_{i}}\right)^{2}\right].$ (35a)
$m_{\tau}(t)=\left(1-\frac{t}{\tau}\right)m_{i}+\frac{t}{\tau}m_{f}$ (35b)
$\sqrt{C_{\tau}(t)}=\left(1-\frac{t}{\tau}\right)\sqrt{C_{i}}+\frac{t}{\tau}\sqrt{C_{f}}$ (35c)
Equation˜35a is the $\mathcal{W}_{2}$ Wasserstein distance between the initial and final Gaussian probability distributions Ref. [39, 40, 44] with the corresponding optimal protocols for the mean and covariance eqs.˜35b and 35c mentioned in Ref.[39].
@@
$\Sigma_{\tau}^{*}=\frac{1}{2(2+\tau)}\left[(m_{f}-m_{i})^{2}+\left(\sqrt{C_{f}}-\sqrt{C_{i}}\right)^{2}\right].$ (36a)
$m_{\tau}(t)=\left(\frac{1+\tau}{2+\tau}-\frac{t}{2+\tau}\right)m_{i}+\left(\frac{1}{2+\tau}+\frac{t}{2+\tau}\right)m_{f}$ (36b)
$\sqrt{C_{\tau}(t)}=\left(\frac{1+\tau}{2+\tau}-\frac{t}{2+\tau}\right)\sqrt{C_{i}}+\left(\frac{1}{2+\tau}+\frac{t}{2+\tau}\right)\sqrt{C_{f}}$ (36c)
Equation˜36a is the novel formulation of the finite-time Wasserstein distance between the initial and final Gaussian probability distributions; it is less than eq.˜35a due to kinks. The finite-time optimal protocol for the center of the trap eq.˜36b is the same as Ref.[7, 8]. However, the finite-time optimal protocol for the trap stiffness eq.˜36c is different from Refs. [7, 8]. These differences, due to boundary conditions and their physical implications, have previously been detailed in section˜3.6.
@@ 5.4 Optimal control of excess EP
The optimal control formulation has been developed for the affinity, which lies in the transition space. However, as previously shown in eq.˜2, a part of the EPR can be integrated to obtain the excess EP $\Sigma^{ex}=-D_{E}^{KL}$, which is defined in the state space. The integration is implemented: 1) from the transition space to the state space, and 2) in time. This implies that $\Sigma^{ex}$ is the boundary term that does not require knowledge of the transition-space topology. This property of the excess EP has been exploited in the literature so far to formulate the control or geodesic description of cEQ systems [74, 16, 75, 44]. For $\Sigma^{ex}=\sum_{\{i\}}\rho_{i}\log{(\rho_{i}/\rho_{i}^{E})}$, the mass is equal to the stochastic Fisher information $\partial^{2}_{\rho_{i}}\Sigma^{ex}=1/\rho_{i}$ defined with respect to $\rho_{i}$  [^1] Note that we use the negative of the excess EP to ensure the positivity of the mass.. Reorganization of ${\Sigma}_{drv}^{ex}=\sum_{\{i\}}(\partial_{t}\rho_{i})^{2}/2\rho_{i}$ leads to ${\Sigma}_{drv}^{ex}=\sum_{\{i\}}\rho_{i}\left[\partial_{t}\log{(\rho_{i}/\rho_{i}^{E})}\right]^{2}$, the quadratic form of the excess driving EP written in terms of the driving of the excess affinity $A_{i}^{ex}=-\log(\rho_{i}/\rho_{i}^{E})$. Thus, the control of the state is equivalent to the control of the excess affinity. It is equal to the Fisher information defined for the state distribution $\{\rho_{i}\}$. To avoid confusion with the state index, we introduce a short-hand notation $\vec{\rho}(\tau)$ and $\vec{\rho}(0)$ for the final and initial states of the system represented as component-wise vectors.
@@
Equation˜37a is the cEQ slow-driving formulation with a quadratic relation between the thermodynamic length and the excess EP [75, 74]. However, the corresponding optimal slow-driving protocol eq.˜37b is novel.
Furthermore, the solution of the GFTOC problem for excess EP leads to the optimal finite-time excess EP and the finite-time optimal driving protocol, which read:
@@
Equation˜38a reveals that the ‘thermodynamic shock’ lowers the driving cost of excess EP due to kinks in $\sqrt{\vec{\rho}_{\tau}(t)}$. Within infinitesimal time at the initial and final driving times, the state of the system undergoes an instantaneous simultaneous jump. Importantly, eqs.˜37 and 38 are constrained by the conservation law manifold. Therefore, they can be further simplified/sophisticated to obtain eqs.˜37 and 38 as a function of the linearly independent state space whose dimension is reduced exactly by the total number of conservation laws.

### v3

@@ 4.1 Multiple control parameters
where eq.˜28 is represented in Einstein’s summation convention, which signifies the summation over the contracted index. $\vec{\lambda}$ is the vector of control parameters. The Hessian of $\mathcal{L}^{*}$ (mass) gives the metric tensor $g_{ij}$,
@@
The MinAP implies solving a variational optimization problem, $\Sigma_{qs}^{*}=\inf(\Sigma_{drv}^{qs})=\int_{0}^{\tau}dt\>\mathcal{L}_{drv}^{*}$. The geodesic equation for the control parameter that minimizes $\Sigma_{drv}^{qs}$ eq.˜28 reads,
@@
$\begin{split}\Gamma_{jk}^{i}=\frac{1}{2}g^{im}\left(\frac{\partial g_{mj}}{\partial{\lambda}^{k}}+\frac{\partial g_{mk}}{\partial\lambda^{j}}-\frac{\partial g_{jk}}{\partial\lambda^{m}}\right)\end{split}$ (31)
Equations˜28 and 30 are equivalent to eqs.˜5 and 7 for the multi-parameter slow-driving optimal control problem.
@@
Using eq.˜32, the optimal slow driving EP reads:
@@
Here, $||*||_{2}$ is the $L^{2}-norm$ that quantifies the distance in the Riemannian manifold induced by multiple control parameters. The finite-time optimal control framework developed in section˜3 is based on the MinAP combined with the existence of the slow-driven geodesic $\mathcal{G}(\{\lambda\})$ and incorporating the possibility of kinks. Hence, a multi-parameter finite-time optimal control protocol is rather trivial, provided that a slow driving geodesic eq.˜32 exists and is computed. The corresponding finite-time multi-parameter optimal protocol reads,
@@
Equations˜32, 33, 34 and 35 are a generalization of eqs.˜8, 9, 21 and 25, respectively, for the multi-parameter generalized optimal control problem.
@@ 4.2 Numerical computation of geodesic
We have focused on the exact analytical formulation of the GFTOC framework. However, analytically computing the exact geodesic $\mathcal{G}(\{\lambda\})$ is not always feasible for more sophisticated problems. To this end, a numerical framework is required to compute the geodesic/optimal protocols. Here, we highlight two different numerical algorithms that are implemented to calculate the geodesic. The first algorithm is based on the MinAP, where the geodesic $\mathcal{G}(\{\lambda\})$ is obtained by the minimization of the Lagrangian eq.˜28, and is extended to return the finite-time optimal protocol $\mathcal{G}_{\tau}(\{\lambda\})$, it reads,
@@
Input: $A_{i},A_{f},\tau,g_{ij}$, eqs.˜28, 34 and 35
1:Compute the geodesic $\mathcal{G}(A_{\alpha})$ obtained by the recursive minimization of eq.˜28, and that connects $\mathcal{G}(A_{i})$ and $\mathcal{G}(A_{f})$.
2:Calculate the finite-time optimal protocol $\mathcal{G}_{\tau}(t)$ and EP cost $\Sigma_{\tau}^{*}$ using eqs.˜34 and 35, respectively.
@@
Note that, Algorithm 1 is modified to directly compute $\mathcal{G}_{\tau}(\{\lambda\})$ instead of $\mathcal{G}(\{\lambda\})$, if the thermodynamic cost for kinks is assigned under the action minimization problem and the endpoint discontinuity in the protocol space is allowed by relaxing the boundary condition constraint. Algorithm 1 extends to the class of numerical optimization methods developed in Ref.[39, 40, 59, 60, 61, 31, 62, 63, 64, 65, 66] and related to machine learning techniques Ref.[67, 68, 69]. However, none of the previous numerical algorithms incorporates kinks as a result of their failure to assign a thermodynamic cost to kinks. To our knowledge, Algorithm 1 is the novel formulation that numerically computes finite-time optimal protocols with kinks and computes the associated finite-time driving EP cost.
A second alternative algorithm computes the finite-time optimal protocols by numerically solving the geodesic eq.˜30 as a boundary value problem. The algorithm for it reads:
@@
Input: $A_{i},A_{f},\tau,g_{ij},\Gamma_{ij}^{k}$, eq.˜30 and eq.˜34
1:Compute the geodesic $\mathcal{G}(A_{\alpha})$ by numerically solving eq.˜30 under the boundary value constraint $A_{i}$ and $A_{f}$ at $t=0$ and $t=\tau$.
2:Calculate the finite-time optimal protocol $\mathcal{G}_{\tau}(t)$ and EP cost $\Sigma_{\tau}^{*}$ using eqs.˜34 and 35, respectively.
@@
In contrast to Algorithm 1, Algorithm 2 can only compute $\mathcal{G}(A_{\alpha})$, and requires eq.˜34 to compute $\mathcal{G}_{\tau}(A_{\alpha})$. The numerical approach elaborated in this section broadens the applicability of the GFTOC when an exact analytical solution of the geodesic is not available.
@@ 5.2 Stochastic particle in a harmonic trap and
A stochastic particle in a harmonic trap potential models an important equivalence class that describes a large class of physical systems, and hence has been the testbed of novel theoretical developments in ST. The stochastic particle satisfies the Boltzmann distribution, $P\propto e^{-k(x-m)^{2}}$, with the trap stiffness $k$ and center $m$. In this case, these two relevant control parameters fully capture the statistical properties of the system. We consider a finite-time optimal control problem for the time-dependent change of the trap stiffness $k(t)$ and center $m(t)$ [10, 24]. The inverse of the trap stiffness quantifies the covariance $(C)$ of the Gaussian distribution for the particle position, $C\propto 1/k$ [10]. Hence, by definition, the finite-time optimal control formulation for changing the centre and stiffness of the harmonic trap is equivalent to the optimal transport theory to change the mean and covariance of the Gaussian distribution for the position of the particles [49] or the application of thermodynamic geometry to the harmonic trap [24]. However, the results obtained from different methods do not match [10, 24, 49, 58]. We resolve these discrepancies using the GFTOC framework.
We consider an optimal control problem for changing the trap stiffness and center from $\{m_{i},k_{i}\}$ to $\{m_{f},k_{f}\}$. The mass for the center and the stiffness of the trap are $g_{mm}=1$ and $g_{kk}=1/k^{3}$ [24]. This implies the geodesic function $\mathcal{G}(m)=m$ and $\mathcal{G}(k)=k^{-1/2}$. Using $C\propto 1/k$, we recover the geodesic function $\mathcal{G}(C)=C^{1/2}$ for the covariance. Hence, the minimum slow driving EP and the corresponding optimal protocols for the mean and variance read:
$\Sigma_{qs}^{*}=\frac{1}{2\tau}\left[(m_{f}-m_{i})^{2}+\left(\sqrt{C_{f}}-\sqrt{C_{i}}\right)^{2}\right],$ (38a)
$m(t)=\left(1-\frac{t}{\tau}\right)m_{i}+\frac{t}{\tau}m_{f},$ (38b)
$\sqrt{C(t)}=\left(1-\frac{t}{\tau}\right)\sqrt{C_{i}}+\frac{t}{\tau}\sqrt{C_{f}}.$ (38c)
Equation˜38a is the square of the $\mathcal{W}_{2}$ Wasserstein distance between the initial and final Gaussian probability distributions divided by $2\tau$ in Refs. [49, 50, 54] with the corresponding optimal protocols for the mean and covariance eqs.˜38b and 38c mentioned in Ref.[49].
@@
$\Sigma_{\tau}^{*}=\frac{1}{2(2+\tau)}\left[(m_{f}-m_{i})^{2}+\left(\sqrt{C_{f}}-\sqrt{C_{i}}\right)^{2}\right],$ (39a)
$m_{\tau}(t)=\left(\frac{1+\tau}{2+\tau}-\frac{t}{2+\tau}\right)m_{i}+\left(\frac{1}{2+\tau}+\frac{t}{2+\tau}\right)m_{f},$ (39b)
$\sqrt{C_{\tau}(t)}=\left(\frac{1+\tau}{2+\tau}-\frac{t}{2+\tau}\right)\sqrt{C_{i}}+\left(\frac{1}{2+\tau}+\frac{t}{2+\tau}\right)\sqrt{C_{f}}.$ (39c)
Equation˜39a is the novel formulation for finite-time processes, which requires dividing the square of $\mathcal{W}_{2}$ Wasserstein distance between the initial and final Gaussian probability distributions by $2(2+\tau)$; which is less than eq.˜38a due to kinks. The finite-time optimal protocol for the center of the trap eq.˜39b is the same as Ref.[10, 11]. However, the finite-time optimal protocol for the trap stiffness eq.˜39c is different from Refs. [10, 11]. These differences, due to boundary conditions and their physical implications, have previously been detailed in section˜3.5.
@@ 5.3 Optimal control of excess EP
The optimal control formulation has been developed for the affinity, which lies in the transition space. However, as previously shown in eq.˜2, a part of the EPR can be integrated to obtain the excess EP $\Sigma^{ex}=-D_{E}^{KL}$, which is defined in the state space. The integration is implemented: 1) from the transition space to the state space, and 2) in time. This implies that $\Sigma^{ex}$ is the boundary term that does not require knowledge of the transition-space topology. This property of the excess EP has been exploited in the literature so far to formulate the control or geodesic description of cEQ systems [83, 23, 84, 54, 85, 86, 87]. For $\Sigma^{ex}=\sum_{\{i\}}\rho_{i}\log{(\rho_{i}/\rho_{i}^{E})}$, the mass is equal to the stochastic Fisher information $\partial^{2}_{\rho_{i}}\Sigma^{ex}=1/\rho_{i}$ defined with respect to $\rho_{i}$  [^6] Note that we use the negative of the excess EP to ensure the positivity of the mass.. Reorganization of ${\Sigma}_{drv}^{ex}=\sum_{\{i\}}(\partial_{t}\rho_{i})^{2}/2\rho_{i}$ leads to ${\Sigma}_{drv}^{ex}=\sum_{\{i\}}\rho_{i}\left[\partial_{t}\log{(\rho_{i}/\rho_{i}^{E})}\right]^{2}$, the quadratic form of the excess driving EP written in terms of the driving of the excess affinity $A_{i}^{ex}=-\log(\rho_{i}/\rho_{i}^{E})$. Thus, the control of the state is equivalent to the control of the excess affinity. To avoid confusion with the state index, we introduce a short-hand notation $\vec{\rho}(\tau)$ and $\vec{\rho}(0)$ for the final and initial states of the system represented as component-wise vectors.
@@
Equation˜40a is the cEQ slow-driving formulation with a quadratic relation between the thermodynamic length and the excess EP [84, 83]. However, the corresponding optimal slow-driving protocol eq.˜40b is novel.
Furthermore, the solution of the novel GFTOC problem for excess EP leads to the optimal finite-time excess EP and the finite-time optimal driving protocol, which read:
@@
Equation˜41a reveals that the ‘thermodynamic shock’ lowers the driving cost of excess EP due to kinks in $\sqrt{\vec{\rho}_{\tau}(t)}$. Within infinitesimal time at the initial and final driving times, the state of the system undergoes an instantaneous simultaneous jump. Importantly, eqs.˜40 and 41 are constrained by the conservation law manifold. Therefore, they can be further simplified/sophisticated to obtain eqs.˜40 and 41 as a function of the linearly independent state space whose dimension is reduced exactly by the total number of conservation laws.
Since, the choice of reference gauge $E$ itself is flexible and should be defined according to physical/experimental constraints, this formulation extends to any reference gauge, for example, $E=ss$ corresponds to any non-equilibrium steady state with a focus on the optimization of the excess EP.
5.4 Linear optimal control of housekeeping EPR
Linear irreversible thermodynamics utilizes the most fundamental quadratic dissipation function $\mathcal{L}^{*}=kA_{\alpha}^{2}$ for the EPR [89, 90], which is also the regime associated with the linear response theories in non-equilibrium thermodynamics. This leads to $\mathcal{L}_{drv}^{*}=k\dot{A}_{\alpha}^{2}$ with the mass $\partial_{A_{\alpha}}^{2}\mathcal{L}^{*}=2k$. Fixing $k=1/2$ implies a unit mass. Solving $\dot{A}_{\alpha}=v_{qs}$ gives $\mathcal{G}^{lin}(A_{\alpha})=A_{\alpha}$. Thus, the geodesic is a linear interpolation between $A_{\alpha}^{i}$ and $A_{\alpha}^{j}$ in $A_{\alpha}$ space, and the slow driving optimal EP and the corresponding optimal protocol read,
The linear optimal control formulated here is a trivial example of Euclidean geometry with a constant mass. This also highlights how, for the cEQ regimes where housekeeping EPR becomes substantially dominant over free energy and excess EP, a trivial linear interpolation between the control parameters leads to a better optimal control procedure than in sections˜5.1 and 5.3.